A man being asked his age replied: "If you take one year from my present age, the result will be three times my son's now, and three years ago, my age was twice what his will be in 5 years." Find the present age of the man.
Answer by ankor@dixie-net.com(12695)

"If you take one year from my present age, the result will be three times my
son's now, and three years ago, my age was twice what his will be in 5 years."
Find the present age of the man.
:
Let m = present age of the man
let s = present age of the son
:
Write an equation for each statement
:
"If you take one year from my present age, the result will be three times my son's now,"
m - 1 = 3s
m = 3s + 1
:
:"three years ago, my age was twice what his will be in 5 years."
m - 3 = 2(s + 5)
m - 3 = 2s + 10
m = 2s + 10 + 3
m = 2s + 13
:
From the 1st statement: substitute (3s+1) for m
3s + 1 = 2s + 13
3s - 2s = 13 - 1
s = 12 yrs is son's present age
then
m = 3(12) + 1
m = 37 yrs is man's present age
:
:
Check solution in the statement:
:"three years ago, my age was twice what his will be in 5 years."
37-3 = 2(12+5)